# No License

### v3

@@ Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This has prompted several studies to explore their representational capability based on the graph isomorphism task. These works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few study GNNs with uncountable node feature representation. In the paper, a novel perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable. More specifically, the strict injective and metric requirements are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation. As a consequence, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the contextualized transformation of neighborhood feature representations via anisotropic and dynamic message functions is proposed. A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies. Experiments on synthetic and benchmark datasets then demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
@@ 1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others (Dwivedi et al., 2023; Hu et al., 2020). GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018a; b). This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph (Li et al., 2015; Murphy et al., 2019; Xu et al., 2018a; Ying et al., 2018).
@@
Among the most widely used GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2016), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Brody et al., 2021; Veličković et al., 2017), and graph isomorphism network (GIN) (Xu et al., 2018a) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework. These models have gained popularity due to their simplicity and remarkable performance across various applications (Dwivedi et al., 2023; Hsu et al., 2021; Hu et al., 2020; Jiang et al., 2022; Kim & Ye, 2020; Liu et al., 2020). Improvements are also constantly being proposed to achieve state-of-the-art performance (Bodnar et al., 2021; Bouritsas et al., 2022; Ishiguro et al., 2019; Miao et al., 2022; Sun et al., 2020; Wang et al., 2019b; Ying et al., 2021).
@@
Notably, these advances are mainly driven by heuristics and empirical results. Nonetheless, several studies have also begun exploring the representational capability of GNNs (Azizian & Lelarge, 2020; Bodnar et al., 2021; Böker et al., 2024; Corso et al., 2020; Garg et al., 2020; Sato et al., 2021). Most of these works analyzed GNNs in relation to the graph isomorphism task. Xu et al. (2018a) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968). Subsequent works build upon their results by considering extensions to the original 1-WL test. However, these results only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. (2020) proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
@@
This paper presents a novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable. The key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation. This idea is investigated across all levels—node-level, neighborhood-level, and graph-level. From the theoretical results, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) which emphasizes the contextualized transformation of neighborhood feature representations using anisotropic and dynamic message functions is proposed. The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the novelty and advantages of the proposed model. Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN, positioning the proposed model as the best-performing MPNN instance.
@@ 2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
$\bm{H_{u}}$ =$=\left\{\!\!\left\{\bm{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}$
$\bm{a_{u}}$ $=\textsc{AGG}\left(\bm{H_{u}}\right)$ (1)
$\bm{h^{*}_{u}}$ $=\textsc{COMB}\left(\bm{h_{u}},\bm{a_{u}}\right),$
where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset (Xu et al., 2018a) of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
@@
When $\mathcal{H}$ is countable, Xu et al. (2018a) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
$F\left(\bm{H}\right)=\sum_{\bm{h}\in\bm{H}}f\left(\bm{h}\right)$ (2)
is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size in the embedded feature space $\mathcal{S}$. This result forms the theoretical basis of GIN.
@@
$M\left(\bm{H}\right)=\bigoplus_{\bm{h}\in\bm{H}}m\left(\bm{h}\right)$ (3)
produces a unique output for every $\bm{H}$ of bounded size. This finding provides the foundation for the principal neighborhood aggregation (PNA) (Corso et al., 2020). Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
@@ 3 Soft-injective functions
While injective functions and metrics are necessary for tasks requiring strict isomorphism, many practical applications of GNNs often do not require such strict constraints. For instance, in node classification tasks, the model must produce identical outputs for some distinct nodes. Thus, this paper softly relaxes these constraints by employing pseudometrics and soft-injective functions.
@@
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$:
  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
[…]
@@
The Euclidean distance is an example of a pseudometric satisfying Assumption 1. A class of pseudometrics satisfying this assumption is provided below, see Berg et al. (1984) and Schölkopf (2000) for more.
@@
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
$d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)<\varepsilon<d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(a) Input feature space $\mathcal{H}$.
$\varepsilon$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(b) Embedded feature space $\mathcal{S}$.
Figure 1: Pseudometric $d_{u}$ and the corresponding feature map $g_{u}$.
@@
Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ may then be described as soft-injective. [^1] The pseudometric $d$ induces the equivalence class $\left[\bm{h}\right]_{d}\coloneqq\left\{\bm{h^{\prime}}\in\mathcal{H}:d\left(\bm{h},\bm{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\bm{h}\right]_{d}:\bm{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938). Hence, $g$ may be described as soft-injective. Corollary 1 extends this result for multisets.

### v4

# Contextualized Messages Boost Graph Representations

@@ Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This has prompted several studies to explore their representational capability based on the graph isomorphism task. Notably, these works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few study GNNs with uncountable node feature representation. In the paper, a new perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable. Specifically, the injective and metric requirements of previous works are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation. As a consequence, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the contextualized transformation of neighborhood feature representations via anisotropic and dynamic message functions is proposed. Furthermore, a mathematical discussion on the relationship between SIR-GCN and key GNNs in literature is laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies. To close, experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
@@ 1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains, including social networks, molecular graphs, and biological networks, among others (Hu et al., 2020a; Dwivedi et al., 2023). In the literature, GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018; 2019). This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the node feature representations of a graph and create an aggregated representation for the entire graph (Ying et al., 2018; Murphy et al., 2019; Xu et al., 2019).
@@
Among the classical GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2017), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Veličković et al., 2018; Brody et al., 2022), and graph isomorphism network (GIN) (Xu et al., 2019) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework. These models have gained popularity due to their simplicity and remarkable performance across various applications (Hu et al., 2020a; Dwivedi et al., 2023). Improvements of these foundational models are also constantly proposed to achieve state-of-the-art performance (Wang et al., 2019b; Bodnar et al., 2021; Bouritsas et al., 2023).
@@
Notably, advances in GNN have mainly been driven by heuristics and empirical results. Nonetheless, several studies have recently begun exploring the representational capability of GNNs (Garg et al., 2020; Sato et al., 2021; Azizian & Lelarge, 2021; Bodnar et al., 2021; Böker et al., 2023). Most of these works analyzed GNNs in relation to the graph isomorphism task. In particular, Xu et al. (2019) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968). Subsequent works build upon their results by considering extensions to the original 1-WL test. Crucially, the theoretical results of these works only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. (2020) proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
@@
This paper presents a new perspective on the representational capability of GNNs when the space of node feature representation is uncountable. Specifically, the key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation. This framework is comprehensively analyzed across all levels—node-level, neighborhood-level, and graph-level. From the theoretical results, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) which emphasizes the contextualized transformation of neighborhood feature representations using anisotropic and dynamic message functions is proposed. This is further accompanied by a discussion on the mathematical relationship between SIR-GCN and key GNNs in literature to underscore the contribution and distinctive advantages of the proposed model. Finally, experiments on synthetic and benchmark datasets in node and graph property prediction tasks are performed to highlight the expressivity of SIR-GCN, positioning the model as a strong candidate for practical GNN applications.
@@ 2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ be the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\boldsymbol{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
$\boldsymbol{H_{u}}$ :-$\coloneq\left\{\!\!\left\{\boldsymbol{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}:$
$\boldsymbol{a_{u}}$ :-$\coloneq\textsc{AGG}\left(\boldsymbol{H_{u}}\right):$ (1)
$\boldsymbol{h^{*}_{u}}$ :-$\coloneq\textsc{COMB}\left(\boldsymbol{h_{u}},\boldsymbol{a_{u}}\right),:$
where AGG and COMB are some aggregation and combination strategies, respectively, $\boldsymbol{H_{u}}$ is the multiset (Xu et al., 2019) of neighborhood features for node $u$, $\boldsymbol{a_{u}}$ is the aggregated neighborhood feature for node $u$, and $\boldsymbol{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
@@
When $\mathcal{H}$ is countable, Xu et al. (2019) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
$F\left(\boldsymbol{H}\right)\coloneq\sum_{\boldsymbol{h}\in\boldsymbol{H}}f\left(\boldsymbol{h}\right)$ (2)
is injective or unique in the embedded feature space $\mathcal{S}$ for each multiset of neighborhood features $\boldsymbol{H}$ of bounded size. This result forms the theoretical basis of GIN.
@@
$M\left(\boldsymbol{H}\right)\coloneq\bigoplus_{\boldsymbol{h}\in\boldsymbol{H}}m\left(\boldsymbol{h}\right)$ (3)
produces a unique output for every $\boldsymbol{H}$ of bounded size. This finding provides the foundation for the principal neighborhood aggregation (PNA) (Corso et al., 2020). Notably, for this result to hold theoretically, the number of aggregators in $\bigoplus$ must scale with the size of the multiset of neighborhood features $\boldsymbol{H}$ which may be impractical for large and dense graphs.
@@ 3 Soft-injective functions
While injective functions and metrics are necessary for tasks requiring graph isomorphism to ensure a unique mapping in the embedded feature space, many practical applications of GNN often do not require such strict constraints. For instance, in node classification tasks, GNNs must produce identical outputs for some distinct nodes. Thus, motivated by Xu et al. (2019) and Corso et al. (2020), this paper softly relaxes the injective and metric requirements within the MPNN framework by employing pseudometrics and soft-injective functions.
@@
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)},\boldsymbol{h}^{(3)}\in\mathcal{H}$:
  1. 1.
    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(1)}\right)=0$;
  2. 2.
    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=d\left(\boldsymbol{h}^{(2)},\boldsymbol{h}^{(1)}\right)$; and
  3. 3.
    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(3)}\right)\leq d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)+d\left(\boldsymbol{h}^{(2)},\boldsymbol{h}^{(3)}\right)$.
Note that for a metric $d$, the first condition is replaced with $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=0\Longleftrightarrow\boldsymbol{h}^{(1)}=\boldsymbol{h}^{(2)}$, ensuring points in $\mathcal{H}$ are distinguishable and unique with respect to $d$. In contrast, for a pseudometric $d$, $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=0$ does not necessarily imply that $\boldsymbol{h}^{(1)}=\boldsymbol{h}^{(2)}$, relaxing the distinguishability constraint of a metric. The following assumption is then imposed on the pseudometric $d$, leveraging results from kernel theory.
Let $\mathcal{H}$ be a non-empty set. A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)},\ldots,\boldsymbol{h}^{(N)}\in\mathcal{H}$,
$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\boldsymbol{h}^{(i)},\boldsymbol{h}^{(j)}\right)\geq 0,$ (4)
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
@@
The Euclidean distance is an example of a pseudometric satisfying Assumption 1. A class of pseudometrics satisfying this assumption is provided below, see Schölkopf (2000) and Berg et al. (2012) for more.
@@
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)<\varepsilon_{1}.$ (5)
@@
Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\boldsymbol{h_{u}}^{(1)},\boldsymbol{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\boldsymbol{h_{u}}^{(1)}$ and $\boldsymbol{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures that $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ may then be described as soft-injective. [^1] The pseudometric $d$ induces the equivalence class $\left[\boldsymbol{h}\right]_{d}\coloneqq\left\{\boldsymbol{h^{\prime}}\in\mathcal{H}:d\left(\boldsymbol{h},\boldsymbol{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\boldsymbol{h}\right]_{d}:\boldsymbol{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938). Fig. 1 provides an illustration of how the soft-injective feature map $g$ maps distinct elements of $\mathcal{H}$ to the same point in $\mathcal{S}$ since the corresponding pseudometric $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=\left\lVert\left[\boldsymbol{h}^{(1)}\right]^{2}-\left[\boldsymbol{h}^{(2)}\right]^{2}\right\rVert$ deems these points to be similar. Corollary 1 extends this result for multisets.
$\left(\mathcal{H},d\right)$${\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\boldsymbol{h^{(1)}}}$${\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\boldsymbol{h^{(2)}}}$${\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\boldsymbol{h^{(3)}}}$$\left(\mathcal{H},d\right)$$\mathcal{S}$$g(d)$${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\mathcal{H}}$$\mathcal{S}$$g\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\boldsymbol{h^{(1)}}}\right)$$g\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\boldsymbol{h^{(2)}}}\right)$$g\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\boldsymbol{h^{(3)}}}\right)$
Figure 1: A soft-injective feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ corresponding to a pseudometric $d$ on $\mathcal{H}$.